19. c + 2x - 3x² is less than unity for all real values of x. the range of value of c is?


1
Expert's answer
2021-02-28T14:53:16-0500

For quadratic expressions such as y=ax2+ bx+ c,

the maximum or minimum value of y=-(b2 -4ac)/4a, where (b2-4ac) is the discriminant of the expression.

The discriminant of the polynomial is a quantity that determines the properties of the roots.

Thus, for this case, b2-4ac= (2)2- 4(-3)(c) = 4+12c

Since c+2x - 3x2 < 1 for all real values of x,

then -(b2-4ac)/4a <1

-(4+12c)/4(-3) <1

(1/3) + c < 1

then the range of value of c is:

c< 2/3